# Copula-like inference for discrete bivariate distributions with rectangular supports

### v2

@@ 1 Introduction
Let $(X,Y)$ be a discrete random vector where $X$ (resp. $Y$) can take $r$ (resp. $s$) distinct values for some strictly positive integers $r$ and $s$. Without loss of generality and for notational convenience, we shall consider that $(X,Y)$ can take its values in $I_{r,s}=I_{r}\times I_{s}$, where $I_{r}=\{1,\dots,r\}$ and $I_{s}=\{1,\dots,s\}$. It is however important to note that the forthcoming developments will only depend on $r$ and $s$ but not on the sets of “labels” $I_{r}$ and $I_{s}$ which were arbitrarily chosen.
As is common in statistics, the probability mass function (p.m.f.) $p$ of $(X,Y)$ is unknown and we instead have at hand $n$ (not necessarily independent) copies $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ of $(X,Y)$ that can be used to carry out inference on $p$. Classical approaches to model $p$ when $r$ and $s$ are not too large are log-linear models (see, e.g., Agresti, 2013; Kateri, 2014; Rudas, 2018) and association models (see, e.g., Goodman, 1985; Kateri, 2014). We shall explore a different class of (semi-parametric or parametric) models in this work hinging upon the possibility of decomposing the unknown p.m.f. $p$ into its two univariate margins and a bivariate p.m.f. with uniform margins playing the role of a discrete copula. A strong case for such a decomposition was indeed recently made by Geenens (2020) in a nice essay on dependence between discrete random variables. The underlying central ingredient is the concept of $I$-projection (in the sense of Csiszar, 1975) on a Fréchet class of p.m.f.s as such projections turn out to preserve the odds ratio matrix which encodes the dependence between $X$ and $Y$ in the considered discrete setting; see Agresti (2013, Section 2.4), Kateri (2014, p 43), Rudas (2018, p 123) or Geenens (2023, Section 4). In practice, the $I$-projection of a p.m.f. on a Fréchet class can be carried out using the iterative proportional fitting procedure (IPFP), also known as Sinkhorn’s algorithm or matrix scaling in the literature. The latter procedure takes the form of an algorithm (actually of several equivalent algorithms) whose aim is to adjust the elements of an input matrix so that it satisfies specified row and column sums. Since its use by Kruithof (1937) for the calculation of telephone traffic, it turned out to play a major role in a surprisingly large number of different scientific contexts (see, e.g., Idel, 2016, for a nice review).
@@ 2.1 Notation
In the context of this work, the set $\Gamma$ thus equivalently represents the set of all bivariate p.m.f.s with $r\times s$ categories whose univariate margins are strictly positive. As we continue, to distinguish bivariate p.m.f.s from other real matrices, we will always use lowercase letters for the former and uppercase letters for the latter. Also, given a bivariate p.m.f. $x$ in $\Gamma$, its first and second univariate margins will be denoted by $x^{[1]}$ and $x^{[2]}$, respectively. With our notation, we thus have
@@
Next, let $a$ and $b$ be fixed p.m.f.s on $I_{r}$ and $I_{s}$, respectively, such that $a_{i}>0$ for all $i\in I_{r}$ and $b_{j}>0$ for all $j\in I_{s}$ and let
@@ 2.2 $I$-projections on Fréchet classes
where, for any $x\in\Gamma$, $\operatorname{supp}(x)=\{(i,j)\in I_{r,s}:x_{ij}>0\}$, and with the convention that $00\log 0=$ 0.
@@
Let $x\in\Gamma$ and let $\mathcal{S}$ be a closed convex subset of $\Gamma$. Suppose furthermore that there exists a p.m.f. $y\in\mathcal{S}$ such that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(x)}$. Then, there exists a unique $y^{*}\in\mathcal{S}$ such that $y^{*}=\operatorname*{arg\,inf}_{y\in\mathcal{S}}D(y\|x)$. Furthermore, we have that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(y^{*})}$ for every $y\in\mathcal{S}$ such that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(x)}$.
@@ 2.3 The iterative proportional fitting procedure
In principle, the IPFP could be applied to any input matrix $x$ in $\mathbb{R}^{r\times s}$ with nonnegative elements whose row and column sums are strictly positive. As this work is of a probabilistic nature, the IPFP will be described only for input matrices $x\in\Gamma$ that can be interpreted as bivariate p.m.f.s. In this context, the aim of the procedure is to adjust an input bivariate p.m.f. $x\in\Gamma$ so that it has $a$ as first margin and $b$ as second margin.
@@ 2.4 Differentiability results for $I$-projections on Fréchet classes
The following proposition, proven in Appendix A, is notationally complex but hinges on the following simple fact: a bivariate p.m.f. with support $T$ can be expressed in terms of $|B|=|T|-1$ elements and a p.m.f. with support $T$ and univariate margins $a$ and $b$ can be expressed in terms of $|A|<|B|$ elements so that the $I$-projection of a matrix in $\Gamma_{T}$ on $\Gamma_{a,b,T}$ can be regarded as a map from an open subset of $(0,1)^{|B|}$ to $(0,1)^{|A|}$. Looking at the aforementioned $I$-projection in such a way allows us to consider it as a map related to an unconstrained optimization problem and, eventually, to apply the implicit function theorem (see, e.g., Fitzpatrick, 2009, Theorem 17.6, p 450).
@@
Using the results given in the previous section, we shall now state a copula-like decomposition for a bivariate p.m.f. on $I_{r,s}$ with strictly positive univariate margins. Let $u^{[1]}$ (resp. $u^{[2]}$) be the univariate p.m.f. of the uniform distribution on $I_{r}$ (resp. $I_{s}$) and let $\Gamma_{\text{unif}}=\Gamma_{u^{[1]},u^{[2]}}$ be the Fréchet class of all bivariate p.m.f.s on $I_{r,s}$ whose univariate margins are $u^{[1]}$ and $u^{[2]}$, respectively. The $I$-projection of a p.m.f. $x\in\Gamma$ on $\Gamma_{\text{unif}}$, if it exists, will be denoted by
@@
In the same way that a continuous bivariate distribution can be be summarized in terms of dependence by a moment of the underlying copula (such as Spearman’s rho or Kendall’s tau – see, e.g., Hofert et al., 2018, Chapter 2 and the references therein), when the decomposition in (3.2) holds, a bivariate p.m.f. $p$ can be summarized by a moment of its copula p.m.f. $u$. In the spirit of Spearman’s rho, as a possible summary of $u$, Geenens (2020) proposed to consider Pearson’s linear correlation $\mathrm{cor}(U,V)$ when $(U,V)$ has p.m.f. $u$. Because of the analogy with Yule’s colligation coefficient (Yule, 1912) for 2 by 2 contingency tables, Geenens (2020, Section 6.6) suggested to call the resulting quantity Yule’s coefficient. Specifically, let $\Upsilon$ be the map from $\Gamma_{\text{unif}}$ to $[-1,1]$ defined, for any $v\in\Gamma_{\text{unif}}$, by
@@
Many other moments of $u$ could be used. As possible alternatives to $\rho$, we shall consider coefficients based on Goodman’s and Kruskal’s gamma (Goodman and Kruskal, 1954) and Kendall’s tau $b$ (see, e.g., Kendall and Gibbons, 1990) of $(U,V)$ when $(U,V)$ has p.m.f. $u$. The following proposition, proven in Appendix B, gives the expressions of Goodman’s and Kruskal’s gamma and Kendall’s tau $b$ for bivariate p.m.f.s on $I_{r,s}$ with standard uniform margins.
@@
[…]
@@
respectively, where the maps $G$ and $T$ are defined in (3.5) and (3.6), respectively. It is important to keep in mind that these coefficients coincide with Goodman’s and Kruskal’s gamma and Kendall’s tau $b$ of $p$ only when $p$ has standard uniform margins.
@@
With the notation of Section 2.3 and given $x\in\Gamma$ with $\operatorname{supp}{(x)}=I_{r,s}$, Theorem 6.2 of Gietl and Reffel (2017) implies that the IPFP of $x$ depends continuously on $x$ and on the underlying target marginals $a$ and $b$. Under a natural assumption of strict positivity of the parametric model for $u$ (see Section 4.2 below), from Proposition 2.4 and the continuous mapping theorem, this implies that $p^{[\alpha^{[n]},\beta^{[n]},\theta^{[n]}]}=\mathcal{I}_{p^{[1,\alpha^{[n]}]},p^{[2,\beta^{[n]}]}}(u^{[\theta^{[n]}]})$ is a consistent estimator of $p$ if $p^{[1,\alpha^{[n]}]}$, $p^{[2,\beta^{[n]}]}$ and $u^{[\theta^{[n]}]}$ are consistent estimators of $p^{[1]}$, $p^{[2]}$ and $u$, respectively. ∎
@@ 4.1 Nonparametric estimation
In our numerical experiments, we considered two possibilities for $q$: the independence copula p.m.f. $\pi$ and, following the suggestion of a Referee, the p.m.f. $\hat{p}^{[n,1]}\hat{p}^{[n,2],\top}$ which has copula p.m.f. $\pi$ and the same margins as $\hat{p}^{[n]}$ in (4.1). Note that the latter choice is only meaningful when both $\hat{p}^{[n,1]}$ and $\hat{p}^{[n,2]}$ are strictly positive. As in our simulations, when applicable, both choices for $q$ led to very similar results, for simplicity, $q$ will be taken equal to $\pi$ as we continue.
@@ 4.2.2 Maximum pseudo-likelihood estimation
Let =$\mathcal{F}=\{p^{[\delta]}:\delta\in\Delta\}$ be a parametric family of bivariate p.m.f.s, where $\Delta$ is a open subset of $\mathbb{R}^{r}$ for some strictly positive integer $r$. Assume furthermore that $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ is a random sample from $p$ (which implies that $n\mathrm{vec}(\hat{p}^{[n]})$ is a multinomial random vector with parameters $n$ and $\mathrm{vec}(p)$, where $\hat{p}^{[n]}$ is defined in (4.1) and $\mathrm{vec}$ is the operator defined as in (2.12) with $S=I_{r,s}$) and that there exists a $\delta_{0}\in\Delta$ such that $p=p^{[\delta_{0}]}$. From (4.15), we see that the maximum pseudo-likelihood estimator in (4.17) bears a strong resemblance with the estimator $\delta^{[n]}=\operatorname*{arg\,inf}_{\delta\in\Delta}D(p^{[n]}\|p^{[\delta]})$. The latter belongs to the classes of minimum divergence estimators of $\delta_{0}$ studied for instance in Read and Cressie (1988), Morales et al. (1995) or Basu et al. (2011). Because the numbers in $nu^{[n]}$ are not observed counts from $u^{[\theta_{0}]}$, the consistency and the asymptotic normality of (4.17) cannot unfortunately be directly deduced from the asymptotic results stated in the aforementioned references. ∎
@@
Assume that the family $\mathcal{J}$ is identifiable (that is, that $\theta\neq\theta^{\prime}$ implies $u^{[\theta]}\neq u^{[\theta^{\prime}]}$) and that there exists $\lambda\in(0,1)$ such that, for any $\theta\in\Theta$ and $(i,j)\in I_{r,s}$, $u^{[\theta]}_{ij}\geq\lambda$. Then, if $\hat{p}^{[n]}\overset{\mathrm{P}}{\to}p$ in $\mathbb{R}^{r\times s}$, where $\hat{p}^{[n]}$ is defined in (4.1), $\theta^{[n]}\overset{\mathrm{P}}{\to}\theta_{0}$ in $\mathbb{R}^{m}$.
@@
where ˙$\dot{\ell}^{[\theta_{0}]}$ is the $m\times rs$ matrix whose column $i+r(j-1)$, $(i,j)\in I_{r,s}$, is ˙$\dot{\ell}^{[\theta_{0}]}_{ij}$ defined as in (4.21) with $\theta=\theta_{0}$ and $J_{u,p}$ is defined in (4.5). Consequently, when $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ are independent copies of $(X,Y)$, the sequence $\sqrt{n}(\theta^{[n]}-\theta_{0})$ is asymptotically centered normal with covariance matrix

### v4

@@ 1 Introduction
Let $(X,Y)$ be a discrete random vector where $X$ (resp. $Y$) can take $r$ (resp. $s$) distinct values for some strictly positive integers $r$ and $s$. As the approach proposed in this work depends only on the values of the probability mass function (p.m.f.) of $(X,Y)$, without loss of generality, we shall consider, for notational convenience only, that $(X,Y)$ takes its values in $I_{r,s}=I_{r}\times I_{s}$, where $I_{r}=\{1,\dots,r\}$ and $I_{s}=\{1,\dots,s\}$.
As is common in statistics, the p.m.f. $p$ of $(X,Y)$ is unknown and we instead have at hand $n$ (not necessarily independent) copies $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ of $(X,Y)$ that can be used to carry out inference on $p$. Classical approaches to model $p$ when $r$ and $s$ are not too large are log-linear models (see, e.g., Agresti, 2013; Kateri, 2014; Rudas, 2018) and association models (see, e.g., Goodman, 1985; Kateri, 2014). We shall explore a different class of (semi-parametric or parametric) models in this work hinging upon the possibility of decomposing the unknown p.m.f. $p$ into its two univariate margins and a bivariate p.m.f. with uniform margins playing the role of a discrete copula. A strong case for such a decomposition was indeed recently made by Geenens (2020) in a nice essay on dependence between discrete random variables. The underlying central ingredient is the concept of $I$-projection (in the sense of Csiszar, 1975) on a Fréchet class of p.m.f.s as such projections turn out to preserve the odds ratio matrix which encodes the dependence between $X$ and $Y$ in the considered discrete setting; see Agresti (2013, Section 2.4), Kateri (2014, p 43), Rudas (2018, p 123) or Geenens (2023, Section 4). In practice, the $I$-projection of a p.m.f. on a Fréchet class can be carried out using the iterative proportional fitting procedure (IPFP), also known as Sinkhorn’s algorithm or matrix scaling in the literature. The latter procedure takes the form of an algorithm (actually of several equivalent algorithms) whose aim is to adjust the elements of an input matrix so that it satisfies specified row and column sums. Since its use by Kruithof (1937) for the calculation of telephone traffic, it turned out to play a major role in a surprisingly large number of different scientific contexts (see, e.g., Idel, 2016, for a nice review).
@@ 2.1 Notation
In the context of this work, the set $\Gamma$ will equivalently represent the set of all bivariate p.m.f.s on $I_{r,s}$ whose univariate margins are strictly positive. As explained in the introduction, since the forthcoming developments depend only on $r$, $s$ and p.m.f. values, restricting attention to p.m.f.s on $I_{r,s}$ is done without loss of generality. Indeed, any p.m.f. of interest defined on the Cartesian product of two sets of reals of cardinalities $r$ and $s$, respectively, can be “relocated” on $I_{r,s}$.
As we continue, to distinguish bivariate p.m.f.s from other real matrices, we will always use lowercase letters for the former and uppercase letters for the latter. Also, given a bivariate p.m.f. $x$ in $\Gamma$, its first and second univariate margins will be denoted by $x^{[1]}$ and $x^{[2]}$, respectively. With our notation, we thus have
@@
Next, let $a$ and $b$ be fixed univariate p.m.f.s on $I_{r}$ and $I_{s}$, respectively, such that $a_{i}>0$ for all $i\in I_{r}$ and $b_{j}>0$ for all $j\in I_{s}$ and let
@@ 2.2 $I$-projections on Fréchet classes
where, for any $x\in\Gamma$, $\operatorname{supp}(x)=\{(i,j)\in I_{r,s}:x_{ij}>0\}$, and with the convention that $00\log 0=$ 0 and $00\log\frac{0}{0}=$ 0.
@@
Let $x\in\Gamma$ and let $\mathcal{S}$ be a closed convex subset of $\Gamma$. Suppose furthermore that there exists a p.m.f. $y\in\mathcal{S}$ such that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(x)}$. Then, there exists a unique $y^{*}\in\mathcal{S}$ such that $y^{*}=\operatorname*{arg\,inf}_{y\in\mathcal{S}}D(y\|x)$. Moreover, we have that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(y^{*})}$ for every $y\in\mathcal{S}$ such that $\operatorname{supp}{(y)}\subset\operatorname{supp}{(x)}$.
@@ 2.3 The iterative proportional fitting procedure
In principle, the IPFP could be applied to any input matrix in $\mathbb{R}^{r\times s}$ with nonnegative elements whose row and column sums are strictly positive. As this work is of a probabilistic nature, the IPFP will be described only for input matrices $x\in\Gamma$ that can be interpreted as bivariate p.m.f.s. In this context, the aim of the procedure is to adjust an input bivariate p.m.f. $x\in\Gamma$ so that it has $a$ as first margin and $b$ as second margin.
@@ 2.4 Differentiability results for $I$-projections on Fréchet classes
The following proposition, proven in Appendix A, is notationally complex but hinges on the following simple fact: a bivariate p.m.f. with support $T$ can be expressed in terms of $|B|=|T|-1$ elements and a p.m.f. with support $T$ and univariate margins $a$ and $b$ can be expressed in terms of $|A|<|B|$ elements so that the $I$-projection of a p.m.f. in $\Gamma_{T}$ on $\Gamma_{a,b,T}$ can be regarded as a map from an open subset of $(0,1)^{|B|}$ to $(0,1)^{|A|}$. Looking at the aforementioned $I$-projection in such a way allows us to consider it as a map related to an unconstrained optimization problem and, eventually, to apply the implicit function theorem (see, e.g., Fitzpatrick, 2009, Theorem 17.6, p 450).
@@
Using the results given in the previous section, we shall now state a copula-like decomposition for a bivariate p.m.f. on $I_{r,s}$ and strictly positive univariate margins. Let $u^{[1]}$ (resp. $u^{[2]}$) be the univariate p.m.f. of the uniform distribution on $I_{r}$ (resp. $I_{s}$) and let $\Gamma_{\text{unif}}=\Gamma_{u^{[1]},u^{[2]}}$ be the Fréchet class of all bivariate p.m.f.s on $I_{r,s}$ whose univariate margins are $u^{[1]}$ and $u^{[2]}$, respectively. The $I$-projection of a p.m.f. $x\in\Gamma$ on $\Gamma_{\text{unif}}$, if it exists, will be denoted by
@@
In the same way that a continuous bivariate distribution can be summarized in terms of dependence by a moment of the underlying copula (such as Spearman’s rho or Kendall’s tau – see, e.g., Hofert et al., 2018, Chapter 2 and the references therein), when the decomposition in (3.2) holds, a bivariate p.m.f. $p$ can be summarized by a moment of its copula p.m.f. $u$. In the spirit of Spearman’s rho, as a possible summary of $u$, Geenens (2020) proposed to consider Pearson’s linear correlation $\mathrm{cor}(U,V)$ when $(U,V)$ has p.m.f. $u$. Because of the analogy with Yule’s colligation coefficient (Yule, 1912) for 2 by 2 contingency tables, Geenens (2020, Section 6.6) suggested to call the resulting quantity Yule’s coefficient. Specifically, let $\Upsilon$ be the map from $\Gamma_{\text{unif}}$ to $[-1,1]$ defined, for any $v\in\Gamma_{\text{unif}}$, by
@@
Many other moments of $u$ could be used. As possible alternatives to $\rho$, we shall consider coefficients based on Goodman’s and Kruskal’s gamma (Goodman and Kruskal, 1954) and Kendall’s tau $b$ (see, e.g., Kendall and Gibbons, 1990) of $(U,V)$ when $(U,V)$ has p.m.f. $u$. The following proposition, proven in Appendix B, gives the expressions of Goodman’s and Kruskal’s gamma and Kendall’s tau $b$ for bivariate p.m.f.s on $I_{r,s}$ with uniform margins.
@@
For any p.m.f. $v\in\Gamma_{\text{unif}}$, Goodman’s and Kruskal’s gamma of $v$ can be expressed as
@@
respectively, where the maps $G$ and $T$ are defined in (3.5) and (3.6), respectively. It is important to keep in mind that these coefficients coincide with Goodman’s and Kruskal’s gamma and Kendall’s tau $b$ of $p$ only when $p$ has uniform margins.
@@
With the notation of Section 2.3 and given $x\in\Gamma$ with $\operatorname{supp}{(x)}=I_{r,s}$, Theorem 6.2 of Gietl and Reffel (2017) implies that the IPFP of $x$ depends continuously on $x$ and on the underlying target marginal p.m.f.s $a$ and $b$. Under a natural assumption of strict positivity of the parametric model for $u$ (see Section 4.2 below), from Proposition 2.4 and the continuous mapping theorem, this implies that $p^{[\alpha^{[n]},\beta^{[n]},\theta^{[n]}]}=\mathcal{I}_{p^{[1,\alpha^{[n]}]},p^{[2,\beta^{[n]}]}}(u^{[\theta^{[n]}]})$ is a consistent estimator of $p$ if $p^{[1,\alpha^{[n]}]}$, $p^{[2,\beta^{[n]}]}$ and $u^{[\theta^{[n]}]}$ are consistent estimators of $p^{[1]}$, $p^{[2]}$ and $u$, respectively. ∎
@@ 4.1 Nonparametric estimation
In our numerical experiments, we considered two possibilities for $q$: the independence copula p.m.f. $\pi$ and, following the suggestion of a Referee, the p.m.f. $\hat{p}^{[n,1]}\hat{p}^{[n,2],\top}$ which has copula p.m.f. $\pi$ and the same margins as $\hat{p}^{[n]}$ in (4.1). Note that the latter choice is only meaningful when both $\hat{p}^{[n,1]}$ and $\hat{p}^{[n,2]}$ are strictly positive (of course, if this is not the case, a practitioner could decide to reduce the cardinalities $r$ or $s$ of the marginal supports). As in our simulations, when applicable, both choices for $q$ led to very similar results, for simplicity, $q$ will be taken equal to $\pi$ as we continue.
@@ 4.2 Parametric estimation
Of course, this way of proceeding is fully meaningful only if the resulting family $\mathcal{J}$ in (4.11) is identifiable, that is, $u^{[\theta]}\neq u^{[\theta^{\prime}]}$ whenever $\theta\neq\theta^{\prime}$. To check that a family $\mathcal{J}$ is identifiable, it thus suffices to verify that, for any $\theta\neq\theta^{\prime}$, there exists $(i,j)\in I_{r,s}$ such that $u^{[\theta]}_{ij}\neq u^{[\theta^{\prime}]}_{ij}$. Note that the quantity $u^{[\theta]}_{ij}$ in (4.12) is actually the $C_{\theta}$-volume of the rectangle $((i-1)/r,i/r]\times((j-1)/s,j/s]$ (see, e.g., Hofert et al., 2018, Section 2.1). Non-identifiability thus occurs when a change in $\theta$ leaves the $C_{\theta}$-volumes of all the rectangles $((i-1)/r,i/r]\times((j-1)/s,j/s]$, $(i,j)\in I_{r,s}$, unchanged. Since the construction in (4.12) is based on classical copula families (which are identifiable), arguably, non-identifiability of the resulting family $\mathcal{J}$ in (4.11) should be the exception rather than the rule in particular as $r$ and $s$ get larger. For a parametric copula family for which $C_{\theta}$ has an explicit expression, identifiability of the family $\mathcal{J}$ can be checked analytically by replacing $C_{\theta}$ by its expression in (4.12). The latter is for instance done in Geenens (2020, Section 7.1) for the Farlie–Gumbel–Morgenstern family. For parametric copula families for which $C_{\theta}$ is not explicitly available (such as elliptical copula families), identifiability could be checked numerically.
@@ 4.2.2 Maximum pseudo-likelihood estimation
Let =$\mathcal{F}=\{p^{[\delta]}:\delta\in\Delta\}$ be a parametric family of bivariate p.m.f.s, where $\Delta$ is a open subset of $\mathbb{R}^{d}$ for some strictly positive integer $d$. Assume furthermore that $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ is a random sample from $p$ (which implies that $n\mathrm{vec}(\hat{p}^{[n]})$ is a multinomial random vector with parameters $n$ and $\mathrm{vec}(p)$, where $\hat{p}^{[n]}$ is defined in (4.1) and $\mathrm{vec}$ is the operator defined as in (2.12) with $S=I_{r,s}$) and that there exists a $\delta_{0}\in\Delta$ such that $p=p^{[\delta_{0}]}$. From (4.15), we see that the maximum pseudo-likelihood estimator in (4.17) bears a strong resemblance with the estimator $\delta^{[n]}=\operatorname*{arg\,inf}_{\delta\in\Delta}D(p^{[n]}\|p^{[\delta]})$. The latter belongs to the classes of minimum divergence estimators of $\delta_{0}$ studied for instance in Read and Cressie (1988), Morales et al. (1995) or Basu et al. (2011). Because the numbers in $nu^{[n]}$ are not observed counts from $u^{[\theta_{0}]}$, the consistency and the asymptotic normality of (4.17) cannot unfortunately be directly deduced from the asymptotic results stated in the aforementioned references. ∎
@@
Assume that the family $\mathcal{J}$ is identifiable and that there exists $\lambda\in(0,1)$ such that, for any $\theta\in\Theta$ and $(i,j)\in I_{r,s}$, $u^{[\theta]}_{ij}\geq\lambda$. Then, if $\hat{p}^{[n]}\overset{\mathrm{P}}{\to}p$ in $\mathbb{R}^{r\times s}$, where $\hat{p}^{[n]}$ is defined in (4.1), $\theta^{[n]}\overset{\mathrm{P}}{\to}\theta_{0}$ in $\mathbb{R}^{m}$.
@@
where ˙$\dot{\ell}^{[\theta_{0}]}$ is the $m\times rs$ matrix whose column $i+r(j-1)$, $(i,j)\in I_{r,s}$, ˙$\dot{\ell}^{[\theta_{0}]}_{ij}$ is defined as in (4.21) with $\theta=\theta_{0}$ and $J_{u,p}$ is defined in (4.5). Consequently, when $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})$ are independent copies of $(X,Y)$, the sequence $\sqrt{n}(\theta^{[n]}-\theta_{0})$ is asymptotically centered normal with covariance matrix